OpenStudy (anonymous):

Solve the exponential equation 3^2x=27

ganeshie8 (ganeshie8):

hint : \(27 = 3\times 3 \times 3 = 3^3\)

OpenStudy (anonymous):

x=3/2

OpenStudy (anonymous):

Can you explain it step by step? I'm confused on how you actually do this

OpenStudy (anonymous):

9*3

OpenStudy (anonymous):

see, as 3^2x = 27, we can see that some number when multiplied by two, and used as a power on 3 give you 27. So you can also write it as 3^2x = 3^3 (as 3*3*3 = 27) Now you can simply use the equation 2x = 3, as you cancelled out the base, because the answer would still remain same. So you get x = 3/2